CRY+VIS Interface Meeting
Date: 2022/4/22 13:00-13:20
Participants: Takahashi, Ikeda, Sato, Yasui, Uchiyama, Aso, Akutsu, Ushiba, Yokozawa
- IX: F0 overload.
- IY: Health check. ACC noise. Cleaning of IYV. Go to finalization.
- EY: MN H1 photosensor. TM actuator. Cleaning of EYV. Go to finalization.
- Type-B: Health check in SRM. Calibration of geophone in SR2. Stand-alone SM driver for SR3.
- Type-Bp: Traverser limit SW.
- Type-C: OFI removal.
- Replacement of flip cables: Cables for TMS were ordered.
- TM height setting: ITMY.
- Oplev: Windshield in the next.
- ACC: Investigating the noise source.
- Commissioning in air:
- Commissioning in vacuum: Alignment of IO and X-arm.
- Acceptance check:
- PF actuator: At the in-vacuum commissioning.
- New vacuum pumps for V-chambers:
- Heater control: Test is pending.
Date: 2022/4/22 14:00-14:10
Participants: Takahashi, Fabian, Uchiyama, Tanaka, Ushiba, Aso, Washimi
- Cleaning of IYV and EYV.
- Takahashi-san compared SR2 IP TF measurements with those of SRM, and found some discrepancies. The problem was likely produced by the geophone integrator being off. In SR2 it has to be turned on by hand. I measured the same TFs this week and they are fine.
- With the aim of understanding Terrece's work, in SR2 I tried to do the inter calibration between geophoes and LVDT. At low frequencies the inter calibration seems be good already, but around 1 Hz, where the resonant frequency of the geophone is, there is a deformation of the TF. I've been looking into this, but I don't have anything relevant to report yet.
- As a result of discussion with the ATC, we add the limit switches to also the stages without the stepper motors. There are some items to be ordered; Jigs (bracket), Touch sensors, PLC board, and In-vac cables.
- Type-A team: Finalization in IY/EY.
- Type-B team:
(Date in Kamioka)
The week from 4/25
- Takahashi: 4/25-28
- Sato: 4/25-28
- Hirata: 4/25-28
The week from 5/2